# Weighted Pseudorandom Generators for Read-Once Branching Programs via Weighted Pseudorandom Reductions

### v2

@@ 3 WPRG for Standard ROBPs
$\left|\mathbb{E}f(U)-\frac{1}{2^{d}}\sum_{i\in 2^{d}}w^{(i)}\mathbb{E}\left[f(\mathsf{R}^{(i)}(U))\right]\right|<\varepsilon,$
@@
Alphabet Reduction: In an alphabet reduction $(\mathsf{R},w)$, each $\mathsf{R}^{(i)}$ is a ($(\{0,1\}^{s^{\prime}})^{n}\to(\{0,1\}^{s})^{n}$ function that maps $n$ short symbols to $n$ long symbols. For any $f\in\mathcal{B}_{(n,s,w)}$, $f\circ\mathsf{R}^{(i)}$ could be computed in $\mathcal{B}_{(n,s^{\prime},w)}$, where the length and width are unchanged but the alphabet log-size is reduced.
Length Reduction: In a length reduction $(\mathsf{R},w)$, each $\mathsf{R}^{(i)}$ is a ($(\{0,1\}^{s^{\prime}})^{k}\to(\{0,1\}^{s})^{n}$ function that maps $k$ long symbols to $n\gg k$ short symbols. For any $f\in\mathcal{B}_{(n,s,w)}$, $f\circ\mathsf{R}^{(i)}$ could be computed in $\mathcal{B}_{(k,s^{\prime},w)}$, where the length is reduced from $n$ to $k$ but the alphabet log-size is increased to $s^{\prime}$.
For the rest of this section, we are going to show the following lemma. The general strategy is to apply the two reductions alternately to reduce the length and maintain the width and alphabet log-size.
@@
For every $\varepsilon\geq 1/\operatorname{\mathrm{poly}}(w)$, there exists an explicit $\varepsilon$-WPRG for $\mathcal{B}_{(n,s,w)}$ with seed length $O\left(s+\frac{\log n\log(nw)}{\max\left\{1,\log\log w-\log\log n\right\}}+\log w(\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon})\right)$ and weight $(8n)^{2\sqrt{\frac{\log w}{\log n}}}$.
We note that for smaller target errors, one can further use the sampler trick [Hoz21] to reduce the error from $1/\operatorname{\mathrm{poly}}(w)$ to an arbitrary $\varepsilon>0$, attaining the following theorem.
@@
$O\left(s+\frac{\log n\log(nw)}{\max\left\{1,\log\log w-\log\log n\right\}}+\log w\left(\log\log\log w-\log\log\max\left\{2,\frac{\log w}{\log n/\varepsilon}\right\}\right)+\log(1/\varepsilon)\right)$
@@ 3.1 Alphabet Reduction
In this section, we will provide a construction of the alphabet reduction. The alphabet reduction reduces the alphabet log-size $s$ to a much smaller $O(\log w)$ and keeps the length $n$ and width $w$ unchanged. The alphabet reduction is achieved by using the Nisan-Zuckerman (NZ) PRG, we start with the construction of the PRG and its main ingredient, extractors.
@@
The original proof of [NZ96] also works for large alphabet. We include a proof in Appendix A for completeness.
@@
Let $\mathrm{E\textsc{xt}}:\{0,1\}^{3s}\times\{0,1\}^{d}\rightarrow\{0,1\}^{s}$ be a $(2s,\frac{\varepsilon}{3n})$-extractor with seed length $d=C_{GUV}\cdot\log\frac{ns}{\varepsilon}$ from Theorem 3.4. Let $\mathsf{NZ}$ be defined as in Lemma 3.5. We define the reduction $(\mathsf{R},w)$:
@@
$w_{x}:=1.$
Now consider that we fix $x$ and let $y_{i}$’s be free. We need to prove that the reduction $(\mathsf{R},w)$ is a $(3s,1,\varepsilon)$-weighted pseudorandom reduction.
@@
Therefore, $(\mathsf{R},w)$ approximates $f$ with error at most $\varepsilon$.
@@ 3.2 Length Reduction framework from Richardson Iteration
A length reduction reduces the length $n$ to a much smaller $k$ but may increase the alphabet log-size. In this subsection, we construct a length reduction using the error reduction given by [AKM+20, CDR+21, PV21], and some explicit PRGs against ROBPs.
@@
To prove Lemma 3.7, we need the following result.
@@
Let $k,K,n_{i,j},\sigma_{i}$ be as in Theorem 3.8, Let $\mathsf{PRG}$ be a $\varepsilon/(n+1)^{2}$-PRG for $\mathcal{B}_{(n,s,w)}$ with seed length $d$ in the assumption. We define the reduction $(\mathsf{R},w)$ as follows:
@@
We will show that $(\mathsf{R},w)$ is a $(\log K,K,\varepsilon^{k}\cdot(n+1))$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(k,d,w)}$.
Let $f\in\mathcal{B}_{(n,s,w)}$. Recall that $f^{[i,j]}(x)$ denotes the transition matrix of $f$ from layer $i$ to layer $j$ with input $x=(x_{1},\ldots,x_{j-i})$. Define $A_{i}=\mathbb{E}_{x\in\{0,1\}^{s}}f^{[i-1,i]}[x]$ and $B_{i,j}=\mathbb{E}_{x\in\{0,1\}^{d}}f^{[i,j]}[\mathsf{PRG}(x)_{j-i}]$. By the definition of the PRG, we have $\|B_{i,j}-A_{i+1}\ldots A_{j}\|_{1}\leq\varepsilon/(n+1)$.
@@
$\left|\mathbb{E}f-\frac{1}{K}\sum_{i=1}^{K}w(i)\mathbb{E}_{X}f(\mathsf{R}_{i}(X))\right|$
$=$ $\left\lVert\mathbb{E}f^{[0,n]}-\sum_{i=1}^{K}\sigma_{i}\mathbb{E}f^{[0,n]}(\mathsf{PRG}(X_{1})_{n_{i,1}},\mathsf{PRG}(X_{2})_{n_{i,2}-n_{i,1}},\ldots,\mathsf{PRG}(X_{k})_{n_{i,k}-n_{i,k-1}})\right\rVert$
@@
$\cdots\mathbb{E}f^{[n_{i,k-1},n_{i,k}]}(\mathsf{PRG}(X_{k})_{n_{i,k}-n_{i,k-1}})\right\rVert$
$=$ $\left\lVert A_{1}\ldots A_{n}-\sum_{i=1}^{K}\sigma_{i}B_{0,n_{i,1}}B_{n_{i,1},n_{i,2}}\ldots B_{n_{i,k-1},n_{i,k}}\right\rVert$
@@ 3.4 Combining the reductions
For all integer $n,s,w$, for all $\varepsilon\geq 1/\operatorname{\mathrm{poly}}(w)$, there exists an explicit
$(O\left(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon})\right),(8n)^{2\sqrt{\frac{\log w}{\log n}}},\varepsilon)$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{\left(O(1),O(\log w),w\right)}$.
@@
Let $(\mathsf{R}^{(1)},w^{(1)})$ be a $(d_{1},W_{1},\varepsilon/4)$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(n_{1},s_{1},w)}$ given by Lemma 3.11, where
@@
  • $W_{1}=(8n)^{\sqrt{\frac{\log w}{\log n}}+1}$.
Let $(\mathsf{R}^{(2)},w^{(2)})$ be a $(s_{2},W_{2},\varepsilon/(4W_{1}))$-weighted pseudorandom reduction from $\mathcal{B}_{(n_{1},s_{1},w)}$ to $\mathcal{B}_{(n_{2},s_{2},w)}$ given by Lemma 3.6, where
@@
  • $W_{2}=1$.
Using Lemma 2.2, $(\mathsf{R}^{(1)}\circ\mathsf{R}^{(2)},w^{(1)}\cdot w^{(2)})$ is a $(d_{1}+d_{2},W_{1},\varepsilon/2)$ weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(\log^{1/2}w,2C_{GUV}\log(1/\varepsilon),w)}$. We set $\varepsilon^{\prime}=(\varepsilon/W_{1})^{2}$ and continue the reduction with the following parameters:
@@
$l=\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon}+O(1).$
For $i=3,\ldots,l-1$, let $(\mathsf{R}^{(i)},w^{(i)})$ be a $(d_{i},K_{i},\varepsilon^{\prime})$-weighted pseudorandom reduction from $\mathcal{B}_{(n_{i},2C_{GUV}\log(1/\varepsilon^{\prime}),w)}$ to $\mathcal{B}_{(n_{i+1},C_{large}\log w,w)}$ given by Lemma 3.10, setting constants $c=3$, $C=\frac{2C_{GUV}\log(1/\varepsilon^{\prime})}{\log w}$, $a=\log_{w}(1/\varepsilon^{\prime})$ in Lemma 3.10, where
@@
For $i=3,\ldots,l-2$, let $(\widehat{\mathsf{R}}^{(i)},\widehat{w}^{(i)})$ be a $(\widehat{d}_{i},\widehat{K}_{i},\varepsilon^{\prime})$-weighted pseudorandom reduction from $\mathcal{B}_{(n_{i+1},C_{large}\log w,w)}$ to $\mathcal{B}_{(n_{i+1},2C_{GUV}\log(1/\varepsilon^{\prime}),w)}$ given by Lemma 3.6, where
@@
We invoke Lemma 2.3 to compose the $72l-$ 7 reductions, which gives $(\mathsf{R}^{(3)}\circ\widehat{\mathsf{R}}^{(3)}\circ\mathsf{R}^{(4)}\circ\widehat{\mathsf{R}}^{(4)}\circ\ldots\circ\mathsf{R}^{(l-1)},w^{(3)}\cdot\widehat{w}^{(3)}\cdot w^{(4)}\cdot\widehat{w}^{(4)}\cdot\cdots\cdot w^{(l-1)})$. Also by Lemma 2.3, we know the reduction is a $(d^{*},W^{*},\varepsilon^{*})$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(n_{l},2C_{GUV}\log(1/\varepsilon^{\prime}),w)}$. The latter class is contained in $\mathcal{B}_{(O(1),C_{GUV}\log w,w)}$.The parameters $(d^{*},W^{*},\varepsilon^{*})$ is shown as following:
@@
  • $W^{*}=\prod_{i=3}^{l-1}W_{i}\cdot\widehat{W}_{i}=\exp(\sum_{i=3}^{l-1}n_{i}^{1/2}\log n_{i})\leq\exp(l\log^{1/3}w)<\exp(\log^{1/2}w)$,
  • $\varepsilon^{*}=\sum_{i=3}^{l-1}\varepsilon^{\prime}\cdot\prod_{j=3}^{i-1}W_{j}\cdot\widehat{W}_{j}\leq l\cdot\varepsilon\cdot W<\varepsilon^{\prime}\cdot\exp(\log^{1/2}w)<\varepsilon/(2W_{1})$.
We invoke Lemma 2.2 again and compose the $(d_{1}+d_{2},W_{1},\varepsilon/2)$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(\log^{1/3}w,2C_{GUV}\log(1/\varepsilon),w)}$ with the $(d^{*},W^{*},\varepsilon^{*})$-weighted pseudorandom reduction from $\mathcal{B}_{(\log^{1/3}w,2C_{GUV}\log(1/\varepsilon),w)}$ to $\mathcal{B}_{(O(1),C_{GUV}\log w,w)}$. We get the desired reduction with the following parameters:
  • Seed length: $s_{1}+s_{2}+s^{*}=O(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon})$,
  • Weight: $W_{1}\cdot W^{*}\leq(8n)^{2\sqrt{\frac{\log w}{\log n}}}$,
  • Error: $\varepsilon/2+W_{1}\cdot\varepsilon/(2W_{1})=\varepsilon$.
@@
For all $C>0,\varepsilon>1/\operatorname{\mathrm{poly}}(w)$, there exists an explicit $\varepsilon$-WPRG for $\mathcal{B}_{(n,C\log w,w)}$ with seed length $O\left(\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon})\right)$ and weight $(8n)^{2\sqrt{\frac{\log w}{\log n}}}$.
@@
Let $(\mathsf{R},w)$ be the reduction from Lemma 3.12. We construct the WPRG $(G,w)$ with the same weight function $w$ and $G$ is defined as $G(x,i)=\mathsf{R}(x,i)$. By Lemma 3.12, to fool the original ROBP we only need to provide randomness for the seed of $(\mathsf{R},w)$ and also provide a random string $x$ which can fool an arbitrary ROBP in $\mathcal{B}_{(O(1),C_{GUV}\log w,w)}$. We use true randomness for $x$ and this only takes $O(\log w)$ random bits. Also the randomness used by $(\mathsf{R},w)$ is $O\left(\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon})\right)$. So the total seed length is as stated. The weight also directly follows from Lemma 3.12. ∎

### v3

@@ 3 WPRG for Standard ROBPs
$\left|\mathbb{E}f(U)-\frac{1}{2^{d}}\sum_{i\in 2^{d}}\sigma^{(i)}\mathbb{E}\left[f(\mathsf{R}^{(i)}(U))\right]\right|<\varepsilon,$
@@
Alphabet Reduction: In an alphabet reduction $(\mathsf{R},\sigma)$, each $\mathsf{R}^{(i)}$ is a ($(\{0,1\}^{s^{\prime}})^{n}\to(\{0,1\}^{s})^{n}$ function that maps $n$ short symbols to $n$ long symbols. For any $f\in\mathcal{B}_{(n,s,w)}$, $f\circ\mathsf{R}^{(i)}$ can be computed in $\mathcal{B}_{(n,s^{\prime},w)}$, where the length and width are unchanged but the alphabet bit-length is reduced.
Length Reduction: In a length reduction $(\mathsf{R},\sigma)$, each $\mathsf{R}^{(i)}$ is a ($(\{0,1\}^{s^{\prime}})^{k}\to(\{0,1\}^{s})^{n}$ function that maps $k$ long symbols to $n\gg k$ short symbols. For any $f\in\mathcal{B}_{(n,s,w)}$, $f\circ\mathsf{R}^{(i)}$ can be computed in $\mathcal{B}_{(k,s^{\prime},w)}$, where the length is reduced from $n$ to $k$ but the alphabet bit-length is increased to $s^{\prime}$.
For the rest of this section, we are going to show the following lemma. The general strategy is to apply the two reductions alternately to reduce the length and maintain the width and alphabet bit-length.
@@
For every $\varepsilon\geq 1/\operatorname{\mathrm{poly}}(w)$, there exists an explicit $\varepsilon$-WPRG for $\mathcal{B}_{(n,s,w)}$ with seed length $O\left(s+\frac{\log n\log(nw)}{\max\left\{1,\log\log w-\log\log n\right\}}+\log w(\log\log\min\{n,\log w\}-\log\log\frac{\log w}{\log(n/\varepsilon)})\right)$ and weight $(8n)^{2\sqrt{\frac{\log w}{\log n}}}$.
We note that for smaller target errors, one can further use the sampler trick [Hoz21] with Lemma 3.1 to reduce the error from $1/\operatorname{\mathrm{poly}}(w)$ to an arbitrary $\varepsilon>0$, attaining the following theorem.
@@
$O\left(s+\frac{\log n\log(nw)}{\max\left\{1,\log\frac{\log w}{\log n}\right\}}+\log w\left(\log\log\min\{n,\log w\}-\log\log\max\left\{2,\frac{\log w}{\log\frac{n}{\varepsilon}}\right\}\right)+\log\frac{1}{\varepsilon}\right).$
@@ 3.1 Alphabet Reduction
The alphabet reduction reduces the alphabet bit-length $s$ to a much smaller $O(\log w)$ and keeps the length $n$ and width $w$ unchanged. The alphabet reduction is achieved by using the Nisan-Zuckerman (NZ) PRG, we start with the construction of the PRG and its main ingredient, extractors.
@@
The original proof of [NZ96] only considers the binary case. So we include a proof in Appendix A for completeness.
@@
Let $\mathrm{E\textsc{xt}}:\{0,1\}^{3s}\times\{0,1\}^{d}\rightarrow\{0,1\}^{s}$ be a $(2s,\frac{\varepsilon}{3n})$-extractor with seed length $d=C_{GUV}\cdot\log\frac{ns}{\varepsilon}$ from Theorem 3.4. Let $\mathsf{NZ}$ be defined as in Lemma 3.5. We define the reduction $(\mathsf{R},\sigma)$:
@@
$\sigma_{x}:=1.$
Now consider that we fix $x$ and let $y_{i}$’s be free. We need to prove that the reduction $(\mathsf{R},\sigma)$ is a $(3s,1,\varepsilon)$-weighted pseudorandom reduction.
@@
Therefore, $(\mathsf{R},\sigma)$ approximates $f$ with error at most $\varepsilon$.
@@ 3.2 Length Reduction framework from Richardson Iteration
A length reduction reduces the length $n$ to a much smaller $k$ but may increase the alphabet bit-length. In this subsection, we construct a length reduction using the error reduction given by [AKM+20, CDR+21, PV21].
@@
To prove Lemma 3.7, we need the following error reduction.
@@
Let $k,K,n_{i,j},\sigma^{\prime}_{i}$ be as in Theorem 3.8, Let $\mathsf{PRG}$ be a $\varepsilon/(2(n+1))$-PRG for $\mathcal{B}_{(n,s,w)}$ with seed length $d$ in the assumption. We define the reduction $(\mathsf{R},\sigma)$ as follows:
@@
We will show that $(\mathsf{R},\sigma)$ is a $(\log K,K,\varepsilon^{k}\cdot(n+1))$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(k,d,w)}$.
Let $f\in\mathcal{B}_{(n,s,w)}$. Recall that $f^{[i,j]}(x)$ denotes the transition matrix of $f$ from layer $i$ to layer $j$ with input $x=(x_{1},\ldots,x_{j-i})$. Define $A_{i}=\mathbb{E}_{x\in\{0,1\}^{s}}f^{[i-1,i]}[x]$ and $B_{i,j}=\mathbb{E}_{x\in\{0,1\}^{d}}f^{[i,j]}[\mathsf{PRG}(x)_{j-i}]$. By the definition of the PRG, we have $\|B_{i,j}-A_{i+1}\ldots A_{j}\|_{\infty}\leq\varepsilon/(n+1)$.
@@
$\left|\mathbb{E}f-\frac{1}{K}\sum_{i=1}^{K}\sigma(i)\mathbb{E}_{X}f(\mathsf{R}_{i}(X))\right|$
$=$ $\left\lVert\mathbb{E}f^{[0,n]}-\sum_{i=1}^{K}\sigma_{i}\mathbb{E}f^{[0,n]}(\mathsf{PRG}(X_{1})_{n_{i,1}},\mathsf{PRG}(X_{2})_{n_{i,2}-n_{i,1}},\ldots,\mathsf{PRG}(X_{k})_{n_{i,k}-n_{i,k-1}})\right\rVert_{\infty}$
@@
$\cdots\mathbb{E}f^{[n_{i,k-1},n_{i,k}]}(\mathsf{PRG}(X_{k})_{n_{i,k}-n_{i,k-1}})\right\rVert_{\infty}$
$=$ $\left\lVert A_{1}\ldots A_{n}-\sum_{i=1}^{K}\sigma_{i}B_{0,n_{i,1}}B_{n_{i,1},n_{i,2}}\ldots B_{n_{i,k-1},n_{i,k}}\right\rVert_{\infty}$
@@ 3.4 Combining the reductions
For all integer $n,s,w$, for all $\varepsilon\geq 1/\operatorname{\mathrm{poly}}(w)$, there exists an explicit $\left(O\left(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\min\{n,\log w\}-\log\log\frac{\log w}{\log n/\varepsilon})\right),(8n)^{2\sqrt{\frac{\log w}{\log n}}},\varepsilon\right)$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{\left(O(1),O(\log w),w\right)}$.
@@
Let $(\mathsf{R}^{(1)},\sigma^{(1)})$ be a $(d_{1},K_{1},\varepsilon/4)$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(n_{1},s_{1},w)}$ given by Lemma 3.11, where
@@
  • $K_{1}=(8n)^{\sqrt{\frac{\log w}{\log n}}+1}$.
Let $(\mathsf{R}^{(2)},\sigma^{(2)})$ be a $(s_{2},K_{2},\varepsilon/(4K_{1}))$-weighted pseudorandom reduction from $\mathcal{B}_{(n_{1},s_{1},w)}$ to $\mathcal{B}_{(n_{2},s_{2},w)}$ given by Lemma 3.6, where
@@
  • $K_{2}=1$.
Using Lemma 2.2, $(\mathsf{R}^{(1)}\circ\mathsf{R}^{(2)},\sigma^{(1)}\cdot\sigma^{(2)})$ is a $(d_{1}+d_{2},K_{1},\varepsilon/2)$ weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(\log^{1/2}w,2C_{GUV}\log(1/\varepsilon),w)}$. We set $\varepsilon^{\prime}=(\varepsilon/K_{1})^{2}$ and continue the reduction with the following parameters:
@@
$l=\log\log\min\{n,\log w\}-\log\log\frac{\log w}{\log n/\varepsilon}+O(1).$
For $i=3,\ldots,l-1$, let $(\mathsf{R}^{(i)},\sigma^{(i)})$ be a $(d_{i},K_{i},\varepsilon^{\prime})$-weighted pseudorandom reduction from $\mathcal{B}_{(n_{i},2C_{GUV}\log(1/\varepsilon^{\prime}),w)}$ to $\mathcal{B}_{(n_{i+1},C_{large}\log w,w)}$ given by Lemma 3.10, setting constants $c=3$, $C=\frac{2C_{GUV}\log(1/\varepsilon^{\prime})}{\log w}$, $a=\log_{w}(1/\varepsilon^{\prime})$ in Lemma 3.10, where
@@
For $i=3,\ldots,l-2$, let $(\widehat{\mathsf{R}}^{(i)},\widehat{\sigma}^{(i)})$ be a $(\widehat{d}_{i},\widehat{K}_{i},\varepsilon^{\prime})$-weighted pseudorandom reduction from $\mathcal{B}_{(n_{i+1},C_{large}\log w,w)}$ to $\mathcal{B}_{(n_{i+1},2C_{GUV}\log(1/\varepsilon^{\prime}),w)}$ given by Lemma 3.6, where
@@
We invoke Lemma 2.3 to compose the $72l-$ 7 reductions, which gives $(\mathsf{R}^{(3)}\circ\widehat{\mathsf{R}}^{(3)}\circ\mathsf{R}^{(4)}\circ\widehat{\mathsf{R}}^{(4)}\circ\ldots\circ\mathsf{R}^{(l-1)},\sigma^{(3)}\cdot\widehat{\sigma}^{(3)}\cdot\sigma^{(4)}\cdot\widehat{\sigma}^{(4)}\cdot\cdots\cdot\sigma^{(l-1)})$. Also by Lemma 2.3, we know the reduction is a $(d^{*},K^{*},\varepsilon^{*})$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(n_{l},2C_{GUV}\log(1/\varepsilon^{\prime}),w)}$. The latter class is contained in $\mathcal{B}_{(O(1),C_{GUV}\log w,w)}$.The parameters $(d^{*},K^{*},\varepsilon^{*})$ is shown as following:
@@
  • $K^{*}=\prod_{i=3}^{l-1}K_{i}\cdot\widehat{K}_{i}=\exp(\sum_{i=3}^{l-1}n_{i}^{1/2}\log n_{i})\leq\exp(l\log^{1/3}w)<\exp(\log^{1/2}w)$,
  • $\varepsilon^{*}=\sum_{i=3}^{l-1}\varepsilon^{\prime}\cdot\prod_{j=3}^{i-1}K_{j}\cdot\widehat{K}_{j}\leq l\cdot\varepsilon\cdot K<\varepsilon^{\prime}\cdot\exp(\log^{1/2}w)<\varepsilon/(2K_{1})$.
We invoke Lemma 2.2 again and compose the $(d_{1}+d_{2},K_{1},\varepsilon/2)$-weighted pseudorandom reduction from $\mathcal{B}_{(n,s,w)}$ to $\mathcal{B}_{(\log^{1/3}w,2C_{GUV}\log(1/\varepsilon),w)}$ with the $(d^{*},K^{*},\varepsilon^{*})$-weighted pseudorandom reduction from $\mathcal{B}_{(\log^{1/3}w,2C_{GUV}\log(1/\varepsilon),w)}$ to $\mathcal{B}_{(O(1),C_{GUV}\log w,w)}$. We get the desired reduction with the following parameters:
  • Seed length: $s_{1}+s_{2}+s^{*}=O(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log(n/\varepsilon)})$,
  • Weight: $K_{1}\cdot K^{*}\leq(8n)^{2\sqrt{\frac{\log w}{\log n}}}$,
  • Error: $\varepsilon/2+K_{1}\cdot\varepsilon/(2K_{1})=\varepsilon$.
@@
For all $C>0,\varepsilon>1/\operatorname{\mathrm{poly}}(w)$, there exists an explicit $\varepsilon$-WPRG for $\mathcal{B}_{(n,C\log w,w)}$ with seed length $O\left(\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log(n/\varepsilon)})\right)$ and weight $(8n)^{2\sqrt{\frac{\log w}{\log n}}}$.
@@
Let $(\mathsf{R},\sigma)$ be the reduction from Lemma 3.12. We construct the WPRG $(G,w)$ with the same weight function $w$ and $G$ is defined as $G(x,i)=\mathsf{R}(x,i)$. By Lemma 3.12, to fool the original ROBP we only need to provide randomness for the seed of $(\mathsf{R},\sigma)$ and also provide a random string $x$ which can fool an arbitrary ROBP in $\mathcal{B}_{(O(1),C_{GUV}\log w,w)}$. We use true randomness for $x$ and this only takes $O(\log w)$ random bits. Also the randomness used by $(\mathsf{R},\sigma)$ is $O\left(\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log(n/\varepsilon)})\right)$. So the total seed length is as stated. The weight also directly follows from Lemma 3.12. ∎